OpenStudy (anonymous):

determine the values of alpha [0,2pi] for which sin(alpha)= cos (alpha/2)

OpenStudy (shamim):

sin 2*(alpha/2)=cos(alpha/2) or.2*sin(alpha/2)*cos(alpha/2)=cos(alpha/2) or. cos(alpha/2)(2*sin(alpha/2)-1)=0

OpenStudy (shamim):

now solve it

OpenStudy (anonymous):

sin alpha/2= 1/2 hence alpha= pi/3 thanks a lot

OpenStudy (shamim):

u r most welcome

OpenStudy (shamim):

we should consider another thing is cos alpha/2=0 or. alpha=180